OpenStudy (anonymous):

A dead tree was struck by lightning, causing it to fall over at a point 10 feet up from the base of the tree. If the fallen portion of the tree (the broken branches) measure 22 feet in length, how far is the top of the tree from the base of the tree?

OpenStudy (anonymous):

A dead tree was struck by lightning, causing it to fall over at a point 10 feet up from the base of the tree. If the fallen portion of the tree (the broken branches) measure 22 feet in length, how far is the top of the tree from the base of the tree?

OpenStudy (anonymous):

can anyone help me with this exercise step by step.

OpenStudy (anonymous):

I think you may have to use the Pythagorean Theorem... Do you know what that is?

jimthompson5910 (jim_thompson5910):

Hint:  now use the Pythagorean Theorem a^2 + b^2 = c^2 10^2 + x^2 = 22^2 to solve for x